# Lorentzian path-integral of Robin Universe

### v1

@@ VIII Discussion and outlook
Our investigations show the important non-trivial role played by the Gauss-Bonnet sector of the gravitational action in favouring the initial configurations which lead to the Hartle-Hawking no-boundary Universe. Gauss-Bonnet term arises in the low Energy-effective action of the heterotic string theory Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx with $\alpha>0$. Although it is topological in nature in four spacetime dimensions and is expected to not play any role in the dynamical evolution, but our analysis clearly shows its contribution to the path integral. When Gauss-Bonnet term is not ignored in path-integral studies, then for $\alpha>0$ (which is also the same sign appearing in low energy effective action of the heterotic string theory) it naturally picks and favors initial configurations which correspond to Hartle-Hawking no-boundary Universe. Moreover, Gauss-Bonnet modifications won’t alter the stability analysis done for Einstein-Hilbert gravity DiTucci:2019dji . This is because Gauss-Bonnet in four spacetime dimension is topological. Perturbative analysis like the one done in DiTucci:2019dji is for small perturbations which are expected to not change the background topology, and hence will be unaffected by the Gauss-Bonnet gravity.
@@ Acknowledgements
We are thankful to the organizers of the international conference QG@RRI 2023 for hosting the wonderful event and giving us a chance to present our work in it. We appreciate the warm hospitality and the vibrant platform the conference provided for exciting discussions.

### v2

@@ VIII Discussion and outlook
Our investigations show the important non-trivial role played by the Gauss-Bonnet sector of the gravitational action in favouring the initial configurations which lead to the Hartle-Hawking no-boundary Universe. Gauss-Bonnet term arises in the low Energy-effective action of the heterotic string theory Cheung:2016wjt ; Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx with $\alpha>0$. Although it is topological in nature in four spacetime dimensions and is expected to not play any role in the dynamical evolution, but our analysis clearly shows its contribution to the path integral. When Gauss-Bonnet term is not ignored in path-integral studies, then for $\alpha>0$ (which is also the same sign appearing in low energy effective action of the heterotic string theory) it naturally picks and favors initial configurations which correspond to Hartle-Hawking no-boundary Universe. Moreover, Gauss-Bonnet modifications won’t alter the stability analysis done for Einstein-Hilbert gravity DiTucci:2019dji . This is because Gauss-Bonnet in four spacetime dimension is topological. Perturbative analysis like the one done in DiTucci:2019dji (see also Ailiga:2024nkz ) is for small perturbations which are expected to not change the background topology, and hence will be unaffected by the Gauss-Bonnet gravity.
@@ Acknowledgements
We are thankful to the organizers of the international conference QG@RRI 2023 for hosting the wonderful event and giving us a chance to present our work in it. We appreciate the warm hospitality and the vibrant platform the conference provided for exciting discussions. We also thank the anonymous referees for their helpful suggestions.
Appendix A Computation of Fluctuation Prefactor
In this appendix, we detail the computation of the fluctuation prefactor for all the boundary conditions discussed in the main text, following Ailiga:2023wzl ; doi:10.1142/7305 . Consider a one-dimensional quantum mechanical system described by the action
$S[q(t)]=\int_{0}^{1}\,\frac{m\dot{q}^{2}}{2}-V(q)\,dt.$ (115)
The path integral we wish to compute is when the Dirichlet boundary condition is imposed on both boundaries (DBC–DBC). The path integral is
$\bar{G}[q_{f};q_{i}]_{DD}=\int_{q(0)=q_{i}}^{q(1)=q_{f}}\mathcal{D}q(t)e^{iS[q(t)]}.$ (116)
The above path integral can be computed exactly for a quadratic potential of the form $V(q)=Aq^{2}+Bq+C$, where, $A,B$, and $C$ are $q$-independent constants. For this potential, the exact path integral acquires the form doi:10.1142/7305
$\bar{G}[q_{f};q_{i}]_{DD}=\mathcal{F}_{DD}e^{iS[q_{cl}(t)]}$ (117)
where,
$\mathcal{F}_{DD}=\left[-\frac{1}{2\pi i\hbar}\frac{\partial^{2}S[q_{cl}(t)]}{\partial q_{i}\partial q_{f}}\right]^{1/2}$ (118)
is the Van Vleck-Pauli-Morette fluctuation determinant for the DBC-DBC case, and $S[q_{cl}(t)]$, is the on-shell action corresponding to the classical solution $q_{cl}(t)$, that satisfies the on-shell equation
$m\ddot{q}_{cl}=-V^{\prime\prime}(q_{cl})$ (119)
subject to the Dirichlet-Dirichlet boundary condition.
Utilizing the above result, we proceed to compute the path integral with Neumann boundary condition (NBC) on the initial hypersurface $p(0)=p_{i}$, and Dirichlet boundary condition on the final hypersurface, $q(1)=q_{f}$. The path integral we wish to evaluate is
$\bar{G}[p_{i};q_{f}]_{ND}=\int_{p(0)=p_{i}}^{q(1)=q_{f}}\mathcal{D}q(t)e^{iS[q(t)]}$ (120)
where, $p(t)$ is the canonical momentum associated to $q(t)$. The above integral again can be evaluated exactly using the following relation Ailiga:2023wzl ; doi:10.1142/7305 ,
[…]
Extending the above definitions and relations to the minisuperspace quantum cosmology, eq. (18), and using the relations in eq. (47), the expression for the DBC–DBC path integral reads Feldbrugge:2017kzv
$\begin{split}\bar{G}_{\rm DBC}[q_{f},t=1;q_{i},t=0]=\sqrt{\frac{3i}{4\pi\hbar N_{c}}}\exp\left[\frac{i}{\hbar}\left\{\frac{N_{c}^{3}\Lambda^{2}}{36}+N_{c}\left(-\frac{\Lambda}{2}(q_{i}+q_{f})+3k\right)\right.\right.\\ \left.\left.+\frac{1}{N_{c}}\left(-\frac{3}{4}(q_{f}-q_{i})^{2}\right)\right\}\right],\end{split}$ (122)
where the exponent corresponds to the on-shell action for the DBC-DBC case, and the prefactor is the Van Vleck-Pauli-Morette fluctuation determinant ($\mathcal{F}_{DD}$). Note that the prefactor matches exactly with the expressions obtained in Ailiga:2023wzl . Furthermore, applying the relation, eq. (121, the path integral for Neumann - Dirichlet boundary conditions is given by
$\bar{G}_{\rm NBC}[q_{f},t=1;\pi_{i},t=0]=\exp\biggl{[}\frac{i}{\hbar}\biggl{\{}\frac{\Lambda^{2}N_{c}^{3}}{9}-\frac{\Lambda\pi_{i}N_{c}^{2}}{3}+\left(3k-\Lambda q_{f}+\frac{\pi_{i}^{2}}{3}\right)N_{c}+q_{f}\pi_{i}\biggr{\}}\biggr{]}\,.$ (123)
Here, as in the DBC–DBC case, the prefactor and exponent correspond to the fluctuation determinant and the on-shell action, respectively. It is crucial to note that the prefactor multiplied to the exponent is unity,
$\mathcal{F}_{ND}=1,$ (124)
which is the fluctuation determinant for the Neumann-Dirichlet case. This agrees with the eq. (49) in the main text and also align with the results obtained in Ailiga:2023wzl ; doi:10.1142/7305 .
One may extend the above analysis for the most general linear Robin boundary condition (RBC) on the initial hypersurface and the Dirichlet boundary condition (DBC) on the final hypersurface. The path integral we wish to evaluate is
$\bar{G}[q_{f};P_{i},\beta]_{RD}=\int_{p(0)+\beta q(0)=P_{i}}^{q(1)=q_{f}}\mathcal{D}q(t)e^{iS[q(t)]}.$ (125)
Again, the above path integral can be done exactly using the following relation Ailiga:2023wzl ,
[…]
Therefore, utilizing the result in eq. (122), the exact expression for the RBC-DBC path integral reads
$\begin{split}\bar{G}_{\rm RBC}[q_{f},t=1;P_{i},t=0]=\frac{1}{\sqrt{1+\frac{2}{3}N_{c}\beta}}\exp\left[\frac{1}{18(3+2N_{c}\beta)}\biggl{\{}\beta\Lambda^{2}N_{c}^{4}+6\Lambda^{2}N_{c}^{3}\right.\\ +\left.N_{c}^{2}\{108\beta k-18\Lambda(P_{i}+\beta q_{f})\}+18N_{c}\left\{9k+P_{i}^{2}+q_{f}\left(\beta^{2}q_{f}-3\Lambda\right)\right\}+54P_{i}q_{f}\biggr{\}}\right].\end{split}$ (127)
where the quantity multiplied to the exponent is the fluctuation prefactor,
$\mathcal{F}_{RD}=\frac{1}{\sqrt{1+\frac{2}{3}N_{c}\beta}}.$ (128)
In the $\beta\rightarrow 0$ limit (NBC-DBC), it goes to eq. 124.